OpenStudy (lena772):

g(x)= 5/x f(x)=7+x/2 give the simplified expression of g(f(x))

OpenStudy (the_fizicx99):

Plug in in the x of f(x) into g(x) g(f(x)) = 5/(7 + x/2)

OpenStudy (yttrium):

\[\ g(f(x)) = \frac{ 5 }{ 7+\frac{ x }{ 2 } }\] \[g(f(x) )= \frac{ 5 }{ \frac{ 14+x }{ 2 } }\] \[g(f(x)) = \frac{ 10 }{14+x }\]

OpenStudy (lena772):

is that simplified fizic?

OpenStudy (the_fizicx99):

No lol, you have to do it xD